A roller of mass 300 kg and of radius 50 cm s lying on  horizontal floor is resting against a step of height 20 cm . the minimum force to be applied on the roller passing through its centre to turn the roller on the step?

Grade:9

2 Answers

Roshan Yarlagadda
6 years ago
Mgh=1/2mv2+1/2Iw2
mgh=1/2mv2+1/2mr2v2/r2
mgh=mv2
v=root gh
v=1.414 m/s
again
W=F.v
1/2mv2+1/2mr2v2/r2=F.v
mv2=F.v
F=m.v
F=300*1.414
F=424.2 N
bharat
6 years ago
i think that for a rolling on a rough inclined plane might help u ,since if these are the steps of ladder then we will have to make equations of motion of rotating body with constant acceleration and also force and mg equations.
if it is inclined then :
acceleration at c.m.=(gsin\Theta/1+I/mr^2)
where \Theta is the angle of the inclined plane  and I is moment of inertia .
please let me know if that step is of ladder or is of incline plane, 
i will make sure that u get the answer of the question afterwards